Thursday, April 14, 2011

Day 16, 04/14/2011 -- G.I. Joe kung fu gripper

Today, we again attacked the problem of the HWP not rotating. We tried a few different things, like hitting the suspect gripper with tons of current and tilting the cryostat to different angles. Still, despite our best efforts, we have yet to find a solution. We do feel we have a better idea of which gripper is not working, and there are some things we can try tomorrow when we go in.

Franky has finished setting up the BROs; tomorrow we plan on connecting cables and cycling the fridges so we can start doing detector work.

Pictures:
https://picasaweb.google.com/asad137/EBEXInPalestine20110414

No comments:

Post a Comment